OpenStudy (anonymous):

Evaluate: 4 − 28 ÷ 7 + 3^2 −9 9 5.6 6

OpenStudy (anonymous):

= 4 - 4 + 3^2 = 3^2 = 3 * 3 = 9

OpenStudy (saifoo.khan):

NEVER FORGET: \[\Huge PEDMAS\]

OpenStudy (anonymous):

*PEMDAS

OpenStudy (saifoo.khan):

http://www.purplemath.com/modules/orderops.htm

OpenStudy (saifoo.khan):

It's PEDMAS. :P

OpenStudy (anonymous):

uh, parentheses, exponents, division, multiplication, addition, and subtraction.

OpenStudy (anonymous):

yep, you're right! xD

OpenStudy (saifoo.khan):

Tx. ;)